OpenStudy (anonymous):

i need help please what is the sum of the finite geometric series?? 1)1+2+4+...+128 2)3+6+12+24+48+...+768 thanks :)

OpenStudy (anonymous):

anybody ?? please :(

OpenStudy (displayerror):

The sum of a finite geometric series is given by \[\sum_{i=1}^{n} a_i = a_1 \frac{1-r^n}{1-r}\] Where \[a_1 = \text{The first term of your geometric series}\] \[r = \text{The common ratio of your geometric series}\] \[n = \text{The number of terms \in your geometric series}\] So for the first geometric series, looking at each term: \[1 + 2 + 4 + \ldots\] What do we multiply 1 by to get 2? What do we multiply 2 by to get 4? Notice that it is the same number--that is our common ratio, r.

OpenStudy (anonymous):

thank you very much!! (;

OpenStudy (anonymous):

i did not get it but thanks anyway

OpenStudy (displayerror):

Which part is confusing you? I'll try to explain it better.

OpenStudy (anonymous):

1) \[S=1+2+4+\ldots+64+128\] \[2S=2(1+2+4+\ldots+64+128)=(2+4+8+\ldots+128+256)\] \[2S-S=S=(2+4+8+\ldots+128+256)-(1+2+4+\ldots+64+128)\] \[S=(\cancel{2}+\cancel{4}+\cancel{8}+\ldots+\cancel{128}+256)-(1+\cancel{2}+\cancel{4}+\ldots+\cancel{64}+\cancel{128})\] \[S=256-1=255\] Now try with 2)